Calculate the number of atoms in each of the following (i) 52 moles of Ar (ii) 52 u of He (iii) 52 g of He.

Grade:12

3 Answers

Suraj Prasad IIT Patna
9 years ago
see i mole contains 6.022 x 1023 atoms.
(i) it says 52 moles of Ar, So 52 x 6.022 x 1023 atoms will be there.

(ii) & (iii) 52 g or 52 u He, both contains no. of moles = 52/4 =13

So, atoms present would be 13 x 6.022 x 1023.
Anjali kumari
5 years ago
As In ( 1) the number of atoms=52×6.022×10^23.And in (2) 13 atoms.And in (3) 13×6.022×10^23. Dit adhura dhyey this dehat evn
ankit singh
3 years ago
(i) 52 moles of Ar
we know,
1 mole of Ar = 6.022 × 10²³ atoms
52 mole of Ar = 52 × 6.022 × 10²³ atoms
= 313.144 × 10²³ atoms
= 3.13144 × 10²⁵ atoms

(ii) 52u of He
we know,
mass of 1 atom of He = 4u
e.g., 4u of He = 1 He atom
∴ 1u of He = 1/4 He atom
∴ 52u of He = 52/4 He atoms = 13 He atoms
(iii) 52g of He
we know,
1 mole of He = 4g = 6.022 × 10²³ atoms of He
no of mole of He of 52g He = 52/4 = 13
∴ no of atoms in 13 mole of He = 13 × Avogadro's constant
= 13 × 6.022 × 10²³ atoms
= 78.286 × 10²³ atoms
= 7.8286 × 10²⁴ atoms